Let the function g : Z → Z be defined by

g(x) = 7x + 3.

(i) Show that g is one-to-one.

(ii) Show that g is not onto.


1
Expert's answer
2021-05-02T08:25:06-0400

"g" is one-to-one iff "\forall x, y\in\Z, g(x)=g(y)=>x=y."

Assume "g(x)=g(y)."

Show it must be true that "x=y"


"g(x)=g(y)=>7x+3=7y+3"

"=>7x=7y"

"=>x=y"

Therefore "g" is one-to-one.


ONTO: Given any "y\in \Z," can we find an "x\in \Z" such that "g(x)=y?"

Counter example:

If "y=0," then


"g(x)=7x+3=0"

"7x=-3"

"x=-\dfrac{3}{7}"

But "-\dfrac{3}{7}" is not an integer. Hence there is no integer "x" for "g(x)=0" and so "g"  is not onto.